# Optimizing Anytime Reasoning via Budget Relative Policy Optimization

### v2

@@ 5 Conclusion
The effectiveness of test-time scaling in LLM reasoning is commonly attributed to the generation-verification gap (Xiang et al., 2025), where verifying solutions is substantially easier than generating them. During reasoning, the model engages in an iterative search process, exploring potential solutions until a valid one is found. Once generated, the solution is verified for correctness, and this search-verification loop continues until a confident answer is produced.
@@ Appendix A Implementation Details
[…]
@@ Appendix B Tree-like Generation and Training
During generation, we use the prefix caching feature of vLLM [Kwon et al., 2023] to reuse computations. We sample a complete thinking process $z$ for a question $x$, then split it based on predefined token budgets ($\{i,j,k\}$ in Figure 9). Each partial thinking process is appended with a special end-of-think token (</think>), and the model is prompted to output the answer directly (see Appendix A for more details).
During training, each response is typically concatenated with its corresponding question using FlashAttention [Dao et al., 2022] for speed. However, this introduces significant duplicated computation for tree-like structures, making it impractical due to high computational demands for LLM training. We implement a tree structure attention mask based on FlexAttention [Li et al., 2024b]. As shown in Figure 9, we append all summaries at the end of the thinking process and record their connection positions in a 1D tensor. This tensor is converted to a block mask by FlexAttention, avoiding 2D tensors that can cause out-of-memory issues for long generation lengths.

### v3

@@ 5 Conclusion
The effectiveness of test-time scaling in LLM reasoning is commonly attributed to the generation-verification gap (metacot), where verifying solutions is substantially easier than generating them. During reasoning, the model engages in an iterative search process, exploring potential solutions until a valid one is found. Once generated, the solution is verified for correctness, and this search-verification loop continues until a confident answer is produced.
@@ Appendix A Implementation Details
Figure 9: We decouple the generation of thinking and its summary. Given the question, the model first generates the thinking, which can be stopped by a special token </think> or the budget limit. Then we insert $\text{Final Answer}**$ (and two ellipsis $\cdots$ plus </think> for out of budget cases) to prompt the model to summarize the answer. In training, these inserted tokens will be ignored when calculating the loss.
@@ Appendix B Tree-like Generation and Training
During generation, we use the prefix caching feature of vLLM [vllm] to reuse computations. We sample a complete thinking process $z$ for a question $x$, then split it based on predefined token budgets ($\{i,j,k\}$ in Figure 10). Each partial thinking process is appended with a special end-of-think token (</think>), and the model is prompted to output the answer directly (see Appendix A for more details).
During training, each response is typically concatenated with its corresponding question using FlashAttention [dao2022flashattention] for speed. However, this introduces significant duplicated computation for tree-like structures, making it impractical due to high computational demands for LLM training. We implement a tree structure attention mask based on FlexAttention [li2024flexattention]. As shown in Figure 10, we append all summaries at the end of the thinking process and record their connection positions in a 1D tensor. This tensor is converted to a block mask by FlexAttention, avoiding 2D tensors that can cause out-of-memory issues for long generation lengths.
@@ D.1 Main Results
D.2 Evaluation on 16k Context Length
We also verified the effectiveness of our method on longer context lengths by fine-tuning DeepSeek-R1-Distill-Qwen-1.5B up to a 16k maximum context. We used a uniform prior over eight thinking budgets: {2k, 4k, 6k, 8k, 10k, 12k, 14k, 16k}. As shown in Figure 12, our approach consistently achieves stronger performance in both anytime and standard reasoning.
Figure 12: The training curves for DeepSeek-R1-Distill-Qwen-1.5B with maximum 16k context length.